OpenStudy (anonymous):

factor : 64x^2 +144x + 81

OpenStudy (anonymous):

(8x-9)^2 i think

OpenStudy (anonymous):

delta = (144)^2 - 4(64)(81) x1 = x2 = 144/2x144